OpenStudy (anonymous):

The function for the cost of materials to make a biscuit is f(x) = four-fifths x + 4, where x is the number of biscuits. The function for the selling price of those biscuits is g(f(x)), where g(x) = 4x + 5. Find the selling price of 15 biscuits.

OpenStudy (anonymous):

i got 65

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

\(\large\color{blue}{ f(x)=\frac{4}{5}x+4 }\) \(\large\color{blue}{ f(x) }\) is the cost, for the number of biscuits \(\large\color{blue}{ x }\).

OpenStudy (anonymous):

so f(15) = 4/5 (15) + 4?

OpenStudy (anonymous):

ohh now i think the answer is 69

OpenStudy (solomonzelman):

\(\large\color{blue}{ f(x)=\frac{4}{5}x+4 }\) function for the price you buy at. \(\large\color{blue}{ g(x)=4x+5 }\) function for the price you sell at. so whatever the +4 in the f(x), and the +5 in g(x) .. maybe the delivery.. but that is the delivery/service or whatever per any number of bisuits x in both function.

OpenStudy (solomonzelman):

both function*s.

OpenStudy (solomonzelman):

So the guy who is selling them makes a big profit... not only off each biscuit, but also from the delivery/service or whatever the y-intercepts are in this case.

OpenStudy (anonymous):

so do i have to solve fot the f one first then do the g one?

OpenStudy (anonymous):

and add them together?

OpenStudy (solomonzelman):

selling price of 15 biscuit would just be, \(\large\color{blue}{ g(15)=4(15)+5 }\) because just like for selling x biscuits you get, the \(\large\color{blue}{ g(x) }\), or 4x+5, so will the selling price be when there are 15 biscuits, saying when x=15.

OpenStudy (anonymous):

so g(15) = 65

OpenStudy (solomonzelman):

yes.

OpenStudy (anonymous):

so i was write in the beginning

OpenStudy (solomonzelman):

The profit on the other hand (for any number of biscuits x) would be \(\large\color{blue}{ g(x)-f(x) }\) or, making my notations, \(\large\color{blue}{ H(x) = }\)\(\large\color{blue}{ g(x)-f(x) }\) \(\large\color{blue}{ H(x) =(4x+5)-(\frac{4}{5}x +4)}\) \(\large\color{blue}{ H(x) =4x+5-\frac{4}{5}x -4}\) \(\large\color{blue}{ H(x) =3\frac{1}{5}x+1}\)

OpenStudy (solomonzelman):

so H(x) is the profit function.

OpenStudy (solomonzelman):

yes, and selling price for 15 biscuits is just g(15) as you have successfully found.

OpenStudy (anonymous):

thanks this is so hard i'm good at math but i'm only supposed to be taking algebra 1 so taking algebra 2 puts so much pressure and it's so confusing

OpenStudy (solomonzelman):

you are doing very good. you seem to comprehend every thing that you are learning.

OpenStudy (solomonzelman):

Just concluding that the selling price is g(15), which is 65... so yw.

OpenStudy (anonymous):

:) i have more if you got time ?